# Tuning-Free Alignment of Diffusion Models with Direct Noise Optimization

### v2

@@ Abstract
In this work, we focus on the alignment problem of diffusion models with a continuous reward function, which represents specific objectives for downstream tasks, such as improving human preference. The central goal of the alignment problem is to adjust the distribution learned by diffusion models such that the generated samples maximize the target reward function. We propose a novel alignment approach, named Direct Noise Optimization (DNO), that optimizes the injected noise during the sampling process of diffusion models. By design, DNO is tuning-free and prompt-agnostic, as the alignment occurs in an online fashion during generation. We rigorously study the theoretical properties of DNO and also propose variants to deal with non-differentiable reward functions. Furthermore, we identify that naive implementation of DNO occasionally suffers from the out-of-distribution reward hacking problem, where optimized samples have high rewards but are no longer in the support of the pretrained distribution. To remedy this issue, we leverage classical high-dimensional statistics theory and propose to augment the DNO loss with certain probability regularization. We conduct extensive experiments on several popular reward functions trained on human feedback data and demonstrate that the proposed DNO approach achieves state-of-the-art reward scores as well as high image quality, all within a reasonable time budget for generation.
@@ 1 Introduction
Figure 2: DNO performs alignment without any modification to the diffusion models. $M_{\theta}$ denotes the sampling process of diffusion models, which maps the noise distribution into the sample distribution. Existing methods for aligning diffusion models seek to modify the noise-to-sample mapping $M_{\theta}$ through either fine-tuning network $\theta$ or gradient-based guidance. In contrast, DNO does not change $M_{\theta}$; instead, it directly optimizes the noise vectors for alignment during inference.
@@
Despite the effectiveness of diffusion models in modeling continuous distributions, a prevalent issue of diffusion models is that they are trained on large-scale, unlabeled, and mixed-quality data to capture the data distribution. However, when deploying these generative models for specific tasks, it is no longer suitable to sample from the original learned distribution directly, because this distribution has not been aligned with the task-specific objective. For instance, in image generation, users aim to generate images that are aesthetically pleasing rather than random ones. Hence, it is imperative to tailor the generative models to align with the task-specific objectives before deployment. The topic of aligning generative models has gained significant attention following the success of aligning LLMs with Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Liu et al., 2023a; OpenAI, 2022; Bai et al., 2022). Recently, the alignment problem has also drawn considerable interest in the context of diffusion models, as evidenced by a series of studies such as (Deng et al., 2024; Uehara et al., 2024; Yuan et al., 2024; Song et al., 2023a; Dong et al., 2023; Sun et al., 2023; Zhang et al., 2023; Prabhudesai et al., 2023; Black et al., 2023; Fan et al., 2023).
Alignment Problem for Diffusion Models. Given a diffusion model characterized by parameters $\theta$ and its associated distribution $p_{\theta}(x)$, as well as a reward function $r(x)$ that can assign real-valued scores to generated samples, the central goal of the alignment problem is to adjust the distribution $p_{\theta}(x)$ such that the generated samples maximize the reward from $r(x)$. Such reward models $r(x)$ can be constructed by training a neural network on human feedback data (Kirstain et al., 2023) or by utilizing custom-designed metrics. In this work, we consider the reward models to be continuous but possibly non-differentiable. In the following sections, we will provide a comprehensive review of some well-established methods for aligning diffusion models.
@@ 1.1.1 Online/Offline Reinforcement Learning for Fine-tuning Diffusion Models
Aligning generative models with specific objectives, such as human preferences, has attracted extensive research efforts, following a series of successful attempts that align LLMs with Reinforcement Learning from Human Feedback (RLHF). For diffusion models, developing alignment algorithms has also gained sufficient attention recently, where a majority of recent research has focused on leveraging Reinforcement Learning (RL)-based approaches. A common mathematical formulation in RL-based methods is to maximize the expected reward while ensuring the resulting distribution does not deviate excessively from the original distribution. This can be expressed as the following KL-regularized optimization problem:
@@
Current RL-based methods can be categorized into online and offline methods based on the data used. In the online method, the algorithm has the capacity to query the reward model throughout the entire optimization process. Two notable online RL methods are DDPO (Black et al., 2023) and DPOK (Fan et al., 2023). Alternatively, research has also delved into the offline RL optimization setting, where an explicit reward model is not accessible and only a fixed preference dataset is utilized. Noteworthy works in this category include Diffusion-DPO (Wallace et al., 2023a) and SPIN-Diffusion (Yuan et al., 2024).
@@ 1.1.3 Loss-Guided Diffusion
A recent work focusing on loss-guided diffusion models (LGD) (Song et al., 2023a) also examines the concept of aligning diffusion models with differentiable rewards. Unlike the methods mentioned previously, LGD is Tuning-Free, meaning it does not necessitate modifications to the pretrained model $\theta$. In essence, the core idea of LGD is that, during the sampling process for the ODE (2), it considers a modified version of the ODE by introducing a new term that guides the generation toward areas of higher reward. Specifically, the new ODE is:
@@
where $x_{0}(x_{t})$ denotes the solution of this ODE starting from $x_{t}$. However, the gradient term $\nabla_{x_{t}}r(x_{0}(x_{t}))$ is not readily available during generation. To address this, the authors suggest utilizing Monte Carlo estimation to approximate the gradient. Nevertheless, this estimation tends to be noisy and imprecise, leading to suboptimal performance, particularly with complex reward models.
@@ 1.2 Comparing Existing Methods
Existing works can be generally categorized based on two criteria: whether it requires fine-tuning and whether the reward model needs to be differentiable.
Tuning-based or Tuning-free Methods. All RL-based methods and the direct fine-tuning method are tuning-based, meaning they necessitate adjustments to the network models $\theta$. There are two main disadvantages associated with tuning-based methods. The first is that they require fine-tuning for each new reward model, a process that can consume considerable resources, especially when faced with extensive choices for reward models. The second drawback is that the fine-tuning process typically relies on a limited set of input prompts, which challenges the model to generalize to new and unseen prompts. In contrast, methods such as LGD (Song et al., 2023a) belong to the tuning-free category. The main advantage of this approach is its elimination of the need for fine-tuning, allowing it to be flexibly applied in conjunction with any reward models and input prompts. Further, tuning-free methods require significantly fewer computing resources than tuning-based methods. However, the major drawback of tuning-free methods is the substantial increase in the time required for the generation process. Also, our experiments in Section 5.2 demonstrate that the tuning-free method LGD cannot match the performance of tuning-based methods such as AlignProp (Prabhudesai et al., 2023) and DDPO (Karunratanakul et al., 2023).
Differentiable or Non-Differentiable Rewards. Current RL-based methods can work by utilizing solely the value or preference information of the reward functions, therefore they can still work if the reward function is non-differentiable. In contrast, the existing direct fine-tuning methods and LGD require the reward function to be differentiable. In practice, working with non-differentiable reward functions is important due to their prevalence. This non-differentiable property can arise from the simulation-based procedures used to compute the reward, or the reward function itself may be a black box, so one cannot directly obtain the gradient information.
@@ 1.3 Our Contributions
In this work, we focus on tuning-free alignment of diffusion models, as we believe that flexibility and low computing requirements are critical for a broad range of real-world applications. Our primary goal is to design a tuning-free method that can achieve or outperform tuning-based methods, and is capable of handling both differentiable and non-differentiable objective functions. To this end, we focus on an under-explored technique for achieving tuning-free alignment of diffusion models—Direct Noise Optimization (DNO). Specifically, we make the following contributions:
@@
  • Through the experiments on several popular image reward functions trained on human feedback data, we demonstrate that our proposed method can achieve state-of-the-art scores in comparison to all existing alignment methods, without any fine-tuning on the models.
@@ 2 Direct Noise Optimization for Aligning Diffusion Models
with $z\sim{\mathcal{N}}(0,I)$ as the initial solution. As we will discuss in Section 3, the Gaussian distribution ${\mathcal{N}}(0,I)$ serves as an important prior on the optimization variables $z$. By solving this optimization problem, we can obtain the optimized noise vectors, which are then used to generate refined samples.
We highlight a prominent difference between the DNO in (8) and major existing alignment methods, including LGD (Song et al., 2023a), DDPO (Black et al., 2023), and AlignProp (Prabhudesai et al., 2023). The core idea of these existing alignment methods is to modify the noise-to-sample mapping $M_{\theta}$ into a new mapping (say $M_{\text{new}}$) such that it can map the Gaussian noise distribution into high-reward samples. The tuning-free method LGD modifies $M_{\theta}$ by augmenting the diffusion ODE/SDE with reward gradient, while the tuning-based methods DDPO and AlignProp modify $M_{\theta}$ by tuning the network parameter $\theta$. Due to this observation, we refer to these existing alignment methods as Alignment Methods with Mapping-Optimization. In sharp contrast, DNO does not seek to modify the mapping $M_{\theta}$. Instead, it aims to find the sub-area of the noise distribution that can be mapped to high-reward samples. See Figure 2 for a visual illustration of this point.
@@
In Section 4, we will reveal that, to efficiently handle non-differentiable reward functions, the best practice for obtaining $\hat{g}(z_{\text{old}})$ is to estimate only the gradient $\nabla_{x}r(x)$, while computing the ground-truth gradient for $\nabla_{z}M_{\theta}(z)$.
We provide a visual illustration in Figure 3 to describe the main procedure for DNO using the DDIM sampling procedure detailed in Algorithm 1. As shown, DNO, similar to LGD Song et al. (2023a), is tuning-free as it does not require tuning the network parameter $\theta$. However, it requires more time for generation compared to direct sampling, as the optimization is integrated with the sampling process, meaning the optimization is performed for each new sample generated. Despite this, as we will demonstrate in Section 5, the extra time needed for the DNO approach is an acceptable trade-off for obtaining high-reward samples.
@@
While the concept of DNO has been proposed in previous works, many of its technical details remain insufficiently explored. On one hand, there is a lack of a comprehensive framework concerning the design choices, theoretical understanding, and practical challenges associated with DNO for aligning diffusion models. On the other hand, the field has yet to systematically investigate whether DNO, as a tuning-free method for aligning diffusion models, can achieve competitive performance compared to tuning-based methods. In this work, we aim to conduct a thorough study on DNO.
@@
2.1 Understanding Noise Optimization
Example 1. First, we present a simple example to visualize the process of noise optimization. Specifically, we trained a toy diffusion model for generating uniform distribution on a ring with a radius between $0.8$ and $1.2$, and the initial distribution is visualized in Figure 4(a), where each red point denotes a single sample drawn from the trained diffusion model. We then solve the noise optimization problem (8) with the reward function $r(x)=\sin(4\pi x[1])+\sin(4\pi x[2])-\left((x[1]-1)^{2}-x[2]^{2}\right)/5$, a highly nonconvex function with many local maxima. To perform this optimization, we solve the DNO problem (8) using gradient ascent with the learning rate set to 0.01. In Figures 4(b), 4(c), and 4(d), we visualize the optimized samples after 10, 50, and 100 gradient steps, respectively. We can observe that the distribution of the samples shifts toward a distribution on the local maxima of the reward function.
@@ 2.1 Understanding Noise Optimization
As described in (11), provided that the step size $\ell$ is less than $\frac{2}{L}$, the distribution $p_{t+1}(x)$ is strictly better than the previous distribution $p_{t}(x)$ in terms of expected reward, as long as the second term is non-zero. Based on this result, it is natural to ask, The next question is: When does the distribution stop improving? Namely, when does the second term in (11) become zero. To answer this question, we provide a detailed discussion on this point and the proof for Theorem 1 in Appendix B. As described in (11), provided that the step size $\ell$ is less than $\frac{2}{L}$, the distribution $p_{t+1}(x)$ is strictly better than the previous distribution $p_{t}(x)$ in terms of expected reward, as long as the second term is non-zero. Based on this result, it is natural to ask: When does the distribution stop improving? Namely, when does the second term in (11) become zero. To answer this question, in Appendix B we provide a detailed discussion, together with the proof for Theorem 1.
@@ 3 Out-Of-Distribution Reward-Hacking in Noise Optimization
It has been observed that when aligning generative models (e.g., including autoregressive language models or diffusion models) with reward functions, one can experience the so-called reward-hacking, i.e., the optimized samples yield high rewards but do not possess the desirable properties (Miao et al., 2024; Chen et al., 2024a). Generally, there are two different types of reward-hacking. In the first type, the reward model is undertrained, so the optimized samples score high rewards but remain barely distinguishable from the samples of the pretrained distribution. The second type is related to the generative model used – the optimized samples no longer fall within the support of the pretrained distribution after optimization. We denote this second type of reward-hacking as Out-Of-Distribution (OOD) Reward-Hacking. In this work, we will focus on this second type and reveal that OOD reward-hacking is a common issue in DNO.
We begin by presenting two examples using both a simple diffusion model from the Example 1 described above and the open-source image diffusion model SD v1.5 (Rombach et al., 2022). For the first example, we revisit the pretrained distribution displayed in Figure 4, modifying the reward function to $r(x)=-(x[1]-1.4)^{2}-(x[2]-1.4)^{2}$. Figure 6(a) exhibits the pretrained distribution of the diffusion models, where we note that every sample stays within the support. However, as illustrated in Figure 6(b), after 1000 gradient steps of optimization for the reward function, all the generated samples become out-of-distribution. In our second example, we examine optimization for the "whiteness" reward—specifically, the average value of all pixels in an image—using SD v1.5 as the image diffusion model. We start with the prompt "black duck", with the initial image depicted in Figure 6(c). After 50 gradient steps of optimization for the whiteness reward, it becomes apparent that the generated samples diverge from the original prompt "black duck" and transform into a white duck, as evidenced in Figure 6(d). Ideally, in the absence of reward-hacking, the generated samples should always adhere to the "black duck" prompt while incorporating the overall whiteness in the images.
@@
The Reasons Behind OOD Samples. In this work, one of our key contributions is to identify one critical cause of OOD reward-hacking in noise optimization. That is, the optimized noise vectors stray towards the low-probability regions of the high-dimensional standard Gaussian distribution; in other words, there is an extremely low chance of such noise vectors being sampled from the Gaussian distribution. As diffusion models are originally trained with Gaussian noise, when the noise vectors originate from these low-probability areas—such as vectors comprised entirely of zeros—the neural network within the diffusion models may incur significant approximation errors for these particular inputs. This error, in turn, prompts the generation of out-of-distribution samples. In the subsequent section, we introduce a novel method to measure the extent to which noise is part of the low-probability region by leveraging the classical concentration inequalities for high-dimensional Gaussian distributions.

### v3

# Inference-Time Alignment of Diffusion Models with Direct Noise Optimization

@@ Abstract
In this work, we focus on the alignment problem of diffusion models with a continuous reward function, which represents specific objectives for downstream tasks, such as increasing darkness or improving the aesthetics of images. The central goal of the alignment problem is to adjust the distribution learned by diffusion models such that the generated samples maximize the target reward function. We propose a novel alignment approach, named Direct Noise Optimization (DNO), that optimizes the injected noise during the sampling process of diffusion models. By design, DNO operates at inference-time, and thus is tuning-free and prompt-agnostic, with the alignment occurring in an online fashion during generation. We rigorously study the theoretical properties of DNO and also propose variants to deal with non-differentiable reward functions. Furthermore, we identify that naive implementation of DNO occasionally suffers from the out-of-distribution reward hacking problem, where optimized samples have high rewards but are no longer in the support of the pretrained distribution. To remedy this issue, we leverage classical high-dimensional statistics theory to an effective probability regularization technique. We conduct extensive experiments on several important reward functions and demonstrate that the proposed DNO approach can achieve state-of-the-art reward scores within a reasonable time budget for generation.
@@ 1 Introduction
Despite the effectiveness of diffusion models in modeling continuous distributions, when deploying these generative models for specific tasks, it is not suitable to sample from the original learned distribution directly, because this distribution has not been aligned with the task-specific objective. For instance, in image generation, users may wish to produce images that are aesthetically pleasing rather than mediocre, or generate images with enhanced brightness, darkness, or compressibility. The topic of aligning generative models has gained significant attention following the success of aligning LLMs with Reinforcement Learning from Human Feedback (RLHF) techniques (Ouyang et al., 2022; Liu et al., 2023a; OpenAI, 2022; Bai et al., 2022). Recently, the alignment problem has also drawn considerable interest in the context of diffusion models, as evidenced by a series of studies such as (Deng et al., 2024; Uehara et al., 2024; Yuan et al., 2024; Song et al., 2023a; Dong et al., 2023; Sun et al., 2023; Zhang et al., 2023; Prabhudesai et al., 2023; Black et al., 2023; Fan et al., 2023).
Alignment Problem for Diffusion Models. Given a diffusion model characterized by parameters $\theta$ and its associated distribution $p_{\theta}(x)$, as well as a reward function $r(x)$ that can assign real-valued scores to generated samples, the central goal of the alignment problem is to adjust the distribution $p_{\theta}(x)$ such that the generated samples maximize the reward from $r(x)$. Such reward function $r(x)$ can be constructed by training a neural network on human feedback data (Kirstain et al., 2023) or by utilizing custom-designed metrics like brightness and darkness. In this work, we consider the reward functions to be continuous but possibly non-differentiable. In the following sections, we will provide a comprehensive review of some well-established methods for aligning diffusion models.
@@ 1.1.1 Online/Offline Reinforcement Learning for Fine-tuning Diffusion Models
Aligning generative models with downstream objectives has attracted extensive research efforts, following a series of successful attempts that align LLMs with Reinforcement Learning from Human Feedback (RLHF). For diffusion models, developing alignment algorithms has also gained sufficient attention recently, where a majority of recent research has focused on leveraging Reinforcement Learning (RL)-based approaches. A common mathematical formulation in RL-based methods is to maximize the expected reward while ensuring the resulting distribution does not deviate excessively from the original distribution. This can be expressed as the following KL-regularized optimization problem:
@@
Current RL-based methods can be categorized into online and offline methods based on the data used. In the online method, the algorithm has the capacity to query the reward function throughout the entire optimization process. Two notable online RL methods are DDPO (Black et al., 2023) and DPOK (Fan et al., 2023), which have been shown to improve downstream objectives such as aesthetics and compressibility. Alternatively, research has also delved into the offline RL optimization setting, where an explicit reward function is not accessible and only a fixed preference dataset is utilized. Noteworthy works in this category include Diffusion-DPO (Wallace et al., 2023a) and SPIN-Diffusion (Yuan et al., 2024).
@@ 1.1.3 Loss-Guided Diffusion
A recent work focusing on loss-guided diffusion models (LGD) (Song et al., 2023a) also examines the concept of aligning diffusion models with differentiable rewards. Unlike the methods mentioned previously, LGD is an Inference-Time alignment method, meaning it does not necessitate modifications to the pretrained model $\theta$ and only works by modifying the inference process. In essence, the core idea of LGD is that, during the sampling process for the ODE (2), it considers a modified version of the ODE by introducing a new term that guides the generation toward areas of higher reward. Specifically, the new ODE is:
@@
where $x_{0}(x_{t})$ denotes the solution of this ODE starting from $x_{t}$. However, the gradient term $\nabla_{x_{t}}r(x_{0}(x_{t}))$ is not readily available during generation. To address this, the authors suggest utilizing Monte Carlo estimation to approximate the gradient. Nevertheless, this estimation tends to be noisy and imprecise, leading to suboptimal performance, particularly with complex reward functions.
@@ 1.2 Comparing Existing Methods
Existing works can be generally categorized based on two criteria: whether it requires fine-tuning and whether the reward function needs to be differentiable.
Inference-Time or Tuning-based Methods. All RL-based methods and the direct fine-tuning method are tuning-based, meaning they necessitate adjustments to the network models $\theta$. There are two main disadvantages associated with tuning-based methods. The first is that they require fine-tuning for new reward functions, a process that can consume considerable resources, especially when faced with extensive choices for reward functions. The second drawback is that the fine-tuning process typically relies on a limited set of input prompts, which challenges the model to generalize to new and unseen prompts. In contrast, methods such as LGD (Song et al., 2023a) belong to the inference-time category. The main advantage of this approach is its elimination of the need for fine-tuning, allowing it to be flexibly applied in conjunction with any reward functions and input prompts. Further, inference-time methods require significantly fewer computing resources than tuning-based methods. However, the major drawback of inference-time methods is the substantial increase in the time required for the generation process.
Differentiable or Non-Differentiable Rewards. Current RL-based methods can work by utilizing solely the value or preference information of the reward functions, therefore they can still work even when the reward function is non-differentiable. In contrast, the existing direct fine-tuning methods and LGD require the reward function to be differentiable. In practice, working with non-differentiable reward functions is important due to their prevalence. This non-differentiable property can arise from the simulation-based procedures used to compute the reward, or the reward function itself may be a black box.
@@ 1.3 Our Contributions
In this work, we focus on inference-time alignment of diffusion models, as we believe that flexibility with the choices of the reward functions, generalization on unseen prompts, and low computing requirements are critical for a broad range of real-world applications. Our primary goal is to design an inference-time alignment method that can match the performance of tuning-based methods by incurring a reasonable additional time cost, and is capable of handling both differentiable and non-differentiable objective functions. To this end, we focus on an under-explored technique for achieving inference-time alignment of diffusion models—Direct Noise Optimization (DNO). Specifically, we make the following contributions:
@@
  • Through the experiments on several important image reward functions, we demonstrate that our proposed method can achieve state-of-the-art scores in comparison to existing alignment methods, without any fine-tuning on the parameters of diffusion models.
@@ 2 Direct Noise Optimization for Aligning Diffusion Models
with $z\sim{\mathcal{N}}(0,I)$ as the initial solution. As we will discuss in Section 3, the Gaussian distribution ${\mathcal{N}}(0,I)$ serves as an important prior on the optimization variables $z$. By solving this optimization problem, we can obtain the optimized noise vectors, which are then used to generate high-reward samples.
@@
In Section 4, we provide a dedicated discussion on how to obtain a better estimator for $\hat{g}(z_{\text{old}})$ when the reward function is non-differentiable.
We provide a visual illustration in Figure 2 to describe the main procedure for DNO using the DDIM sampling procedure detailed in Algorithm 1. As shown, DNO, similar to LGD (Song et al., 2023a), operates at inference-time and does not require tuning the network parameter $\theta$. However, it requires more time for generation compared to direct sampling, as the optimization is integrated with the sampling process, meaning the optimization is performed for each new sample generated. Despite this, as we will demonstrate in Section 5, the extra time needed for the DNO approach is a worthwhile trade-off for obtaining high-reward samples.
@@
While similar methods of DNO has appeared in previous works, many of its technical details remain insufficiently explored. On one hand, there is a lack of a comprehensive framework concerning the design choices, theoretical understanding, and practical challenges associated with DNO for aligning diffusion models. On the other hand, the field has yet to systematically investigate whether DNO, as a inference-time method for aligning diffusion models, can achieve competitive performance compared to tuning-based methods. In this work, we aim to conduct a thorough study on DNO.
@@
2.1 Understanding Direct Noise Optimization
Example 1. First, we present a simple example to visualize the process of noise optimization. Specifically, we trained a toy diffusion model for generating uniform distribution on a ring with a radius between $0.8$ and $1.2$, and the initial distribution is visualized in Figure 3(a), where each red point denotes a single sample drawn from the trained diffusion model. We then solve the noise optimization problem (8) with the reward function $r(x)=\sin(4\pi x[1])+\sin(4\pi x[2])-\left((x[1]-1)^{2}-x[2]^{2}\right)/5$, a highly nonconvex function with many local maxima. To perform this optimization, we solve the DNO problem (8) using gradient ascent with the learning rate set to 0.01. In Figures 3(b), 3(c), and 3(d), we visualize the optimized samples after 10, 50, and 100 gradient steps, respectively. We can observe that the distribution of the samples shifts toward a distribution on the local maxima of the reward function.
@@ 2.1 Understanding Direct Noise Optimization
As described in (11), provided that the step size $\ell$ is less than $\frac{2}{L}$, the distribution $p_{t+1}(x)$ is strictly better than the previous distribution $p_{t}(x)$ in terms of expected reward, as long as the second term is non-zero.
Based on this result, it is natural to ask: When does the distribution stop improving? Namely, when does the second term in (11) become zero. We provide a detailed discussion to answer this question and also the proof for Theorem 1 in Appendix B.
@@ 3 Out-Of-Distribution Reward-Hacking in Noise Optimization
It has been observed that when aligning generative models (e.g., including autoregressive language models or diffusion models) with reward functions, one can experience the so-called reward-hacking, i.e., the optimized samples yield high rewards but do not possess the desirable properties (Miao et al., 2024; Chen et al., 2024a). Generally, there are two different types of reward-hacking. In the first type, the reward function admits some shortcuts, so the optimized samples score high rewards but remain barely distinguishable from the samples of the pretrained distribution. The second type is related to the generative model used – the optimized samples no longer fall within the support of the pretrained distribution after optimization. We denote this second type of reward-hacking as Out-Of-Distribution (OOD) Reward-Hacking. In this work, we will focus on this second type and reveal that OOD reward-hacking is a common issue in DNO.
We begin by presenting two examples using both a simple diffusion model from the Example 1 described above and the open-source image diffusion model SD v1.5 (Rombach et al., 2022). For the first example, we revisit the pretrained distribution displayed in Figure 3, modifying the reward function to $r(x)=-(x[1]-1.4)^{2}-(x[2]-1.4)^{2}$. Figure 5(a) exhibits the pretrained distribution of the diffusion models, where we note that every sample stays within the support. However, as illustrated in Figure 5(b), after 1000 gradient steps of optimization for the reward function, all the generated samples become out-of-distribution. In our second example, we examine optimization for the "brightness" reward—specifically, the average value of all pixels in an image—using SD v1.5 as the image diffusion model. We start with the prompt "black duck", with the initial image depicted in Figure 5(c). After 50 gradient steps of optimization for the brightness reward, it becomes apparent that the generated samples diverge from the original prompt "black duck" and transform into a "white duck", as evidenced in Figure 5(d). Ideally, in the absence of reward-hacking, the generated samples should always adhere to the "black duck" prompt while incorporating the overall brightness in the images.
@@
The Reasons Behind OOD Samples. In this work, one of our key contributions is to identify one critical cause of OOD reward-hacking in noise optimization. That is, the optimized noise vectors stray towards the low-probability regions of the high-dimensional standard Gaussian distribution; in other words, there is an extremely low chance of such noise vectors being sampled from the Gaussian distribution. As diffusion models are originally trained with Gaussian noise, when the noise vectors originate from these low-probability areas—such as vectors comprised entirely of zeros—the neural network within the diffusion models may incur significant approximation errors for these particular inputs. This error, in turn, leads to the generation of out-of-distribution samples. In the subsequent section, we introduce a novel method to measure the extent to which noise is part of the low-probability region by leveraging the classical concentration inequalities for high-dimensional Gaussian distributions.